OpenStudy (anonymous):

Determine if triangle RST with coordinates R (3, 4), S (5, 5), and T (6, 1) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific. @phi

jimthompson5910 (jim_thompson5910):

Find the slope of these segments: RS, ST, RT Use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] If the product of any two slopes is -1, then those two segments are perpendicular, which leads to a right angle. If you have a right angle, you have a right triangle.

OpenStudy (anonymous):

Sorry, I just saw this. I'll start on figuring that out right now. :)

jimthompson5910 (jim_thompson5910):

ok tell me what slopes you get

OpenStudy (anonymous):

http://sketchtoy.com/65524992

jimthompson5910 (jim_thompson5910):

You mixed up some of the numbers I think. The first calculation and third calculation is what I'm referring to. The second calculation is correct.

OpenStudy (anonymous):

what did I do incorrectly with those?

jimthompson5910 (jim_thompson5910):

you have the correct slope for ST (middle calculation) so I'm going to do the other ones

jimthompson5910 (jim_thompson5910):

R = (x1,y1) = (3,4) x1 = 3 y1 = 4 S = (x2,y2) = (5,5) x2 = 5 y2 = 5 ------------------------------------ slope of RS \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{5-4}{5-3}\] \[\Large m = \frac{1}{2}\]

jimthompson5910 (jim_thompson5910):

R = (x1,y1) = (3,4) x1 = 3 y1 = 4 T = (x2,y2) = (6,1) x2 = 6 y2 = 1 ------------------------------------ slope of RT \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] \[\Large m = \frac{1-4}{6-3}\] \[\Large m = \frac{-3}{3}\] \[\Large m = -1\]

OpenStudy (anonymous):

Oh, I see now. Thanks! What do I do next?

jimthompson5910 (jim_thompson5910):

Summary of slopes slope of RS = 1/2 = 0.5 slope of ST = -4 slope of RT = -1

jimthompson5910 (jim_thompson5910):

do any of those slopes pair up and multiply to -1? if so, which two slopes do this?

OpenStudy (anonymous):

None of them.

jimthompson5910 (jim_thompson5910):

correct, so you don't have any perpendicular segments that means you don't have any right angles

OpenStudy (anonymous):

How do I show them how to make it a right angle?

jimthompson5910 (jim_thompson5910):

You would have to move one of the points to a place where you can get two slopes to multiply to -1

OpenStudy (anonymous):

Awesome. Thanks!

jimthompson5910 (jim_thompson5910):

For instance, moving point T from (6,1) to (7,1) will make slope of ST = -2 multiply the new slope of ST (-2) with the slope of RS (0.5) to get -2*0.5 = -1. So ST and RS are perpendicular. You have a right angle at point S. This is of course only after you move T to (7,1)